# CleanMel: Mel-Spectrogram Enhancement for Improving Both Speech Quality and ASR

### v1

@@ Abstract
In this work, we propose CleanMel, a single-channel Mel-spectrogram denoising and dereverberation network for improving both speech quality and automatic speech recognition (ASR) performance. The proposed network takes as input the noisy and reverberant microphone recording and predicts the corresponding clean Mel-spectrogram. The enhanced Mel-spectrogram can be either transformed to speech waveform with a neural vocoder or directly used for ASR. The proposed network is composed of interleaved cross-band and narrow-band processing in the Mel-frequency domain, for learning the full-band spectral pattern and the narrow-band properties of signals, respectively. Compared to linear-frequency domain or time-domain speech enhancement, the key advantage of Mel-spectrogram enhancement is that Mel-frequency presents speech in a more compact way and thus is easier to learn, which will benefit both speech quality and ASR. Experimental results on four English and one Chinese datasets demonstrate a significant improvement in both speech quality and ASR performance achieved by the proposed model. Code and audio examples of our model are available online  [^1] https://audio.westlake.edu.cn/Research/CleanMel.html.
@@ I Introduction
THIS work studies single-channel speech enhancement using deep neural networks (DNNs), to improve both speech quality and automatic speech recognition (ASR) performance. A large class of speech enhancement methods employ DNNs to map from noisy and reverberant speech to corresponding clean speech, conducted either in time domain [1, 2] or in time-frequency domain [3, 4, 5]. These methods can efficiently suppress noise, but not necessarily improve ASR performance due to the speech artifacts/distortions caused by speech enhancement networks [6]. In [7], it was found that time-domain enhancement is more ASR-friendly than frequency-domain enhancement. A time domain progressive learning method is proposed in [8], which also shows the superiority of time domain speech enhancement, and the progressive learning mechanism is very effective for robust ASR by mitigating the over suppression of speech. In [9, 10], ASR performance is largely improved by decoupling frontend enhancement and backend recognition. In [10], it is shown that an advanced time-frequency domain network, i.e. CrossNet, can even outperform the time domain ARN network.
@@
Speech enhancement in Mel-frequency domain, or similarly in rectangular bandwidth (ERB) domain, has been developed under various contexts in the literature. Mel-frequency and ERB bands model human speech perception of spectral envelope and signal periodicity, within which speech enhancement is more perceptually and computationally efficient than within linear-frequency domain or time domain. In [11, 12, 13], spectral envelope enhancement is performed within the ERB bands, and then applying pitch filtering [11, 12] or deep filtering [13] to recover the enhanced speech. Sub-band networks [3, 14] and full-band/sub-band fusion networks, i.e. FullSubNet [15, 16], have been recently proposed and achieved outstanding performance. However, separately processing sub-bands in the linear-frequency domain leads to a large computational complexity. To reduce the number of sub-bands and thus the computational complexity, Fast FullSubNet [17] and the work of [18] proposed to perform sub-band processing in the Mel-frequency domain, and then transform back to linear frequency with a joint post-processing network. In [19, 20], speech enhancement is directly conducted in the Mel-frequency domain and then a separate neural vocoder is used to recover speech waveform. These methods improve speech enhancement capability by alleviating the burden of enhancing full-band speech details and also by leveraging the powerful full-band speech recovery capacity of advanced neural vocoder, as a result, achieve higher speech quality compared to their linear-frequency counterparts.
@@
In this work, we propose CleanMel, a single-channel Mel-spectrogram enhancement network for improving both speech quality and ASR performance. Different from the previous works [11, 12, 13, 17] that performing speech enhancement in the ERB or Mel domain, and then applying a joint pitch filtering or deep filtering to obtain the enhanced speech, this work decouples the Mel-spectrogram enhancement and post-processing parts by targeting the enhancement network with clean Mel-spectrogram. The enhanced Mel-spectrogram can be directly used for ASR, or transformed back to waveform with a separate neural vocoder as is done in [19]. Compared to linear-frequency spectrogram or time-domain waveform, Mel-frequency presents speech in a more compact and less-detailed way (but still perceptually efficient) and has a lower feature dimension (number of frequencies) from the perspective of machine learning, which would result in lower the prediction error. This is beneficial for both speech quality improvement and ASR: (i) Neural vocoders have been extensively studied in the field of Text-to-Speech, and are capable of efficiently transforming Mel-spectrogram back to time-domain waveform. Therefore, the low-error property of Mel-spectrogram enhancement can be hopefully maintained by the neural vocoders and higher speech quality can be achieved. (ii) From the perspective of ASR, there seems no need to first recover the less-accurate full-band speech details and then compress to Mel-frequency. A direct and more-accurate Mel-spectrogram estimation would be preferred.
@@
We adapt the network architecture of our previous proposed (online) SpatialNet [21, 22] with some modifications to better accommodate Mel-spectrogram enhancement. SpatialNet is composed of interleaved cross-band and narrow-band blocks originally proposed for processing multichannel STFT frequencies. The narrow-band block processes STFT frequencies independently to learn the spatial information presented in narrow-band (one frequency), such as the convolutive signal propagation and the spatial correlation of noise. And the cross-band block was designed for learning the across-frequency dependencies of narrow-band information. As for single-channel Mel-spectrogram enhancement in this work, the narrow-band block processes Mel frequencies independently to also learn the (single-channel) convolutive signal propagation of target speech, which is crucial not only for conducting dereverberation of target speech but also for discriminating between target speech and interfering signals. The cross-band block is now reinforced and utilized to learn the full-band spectral pattern in the Mel-frequency domain.
@@
In addition, we have studied several critical issues when decoupling the Mel-spectrogram enhancement front-end and ASR/Vocoder back-ends. i) We have systematically studied and compared different learning targets that can be used for Mel-spectrogram enhancement, including logMel mapping, Mel ratio masking and the clipping issue of logMel; ii) We have developed a data normalization scheme to align the signal levels of cascaded front-end and back-end models; iii) We have developed an online neural vocoder to enable online speech enhancement.
@@
Experiments are conducted on five public datasets (four English and one Chinese) for speech denoising and dereverberation individually or jointly. Importantly, we adopt a more realistic evaluation setup: from multiple data sources of clean speech, real-measured room impulse responses (RIRs) and noise signals, we collected and organized a relatively large-scale training set, based on which we train the network for once and directly test it on all the five test sets. Experiments show that the proposed model achieves the state-of-the-art (SOTA) speech enhancement performance in term of speech perceptual quality. Moreover, on top of various pre-trained and advanced ASR models, the proposed model prominently improves the ASR performance on all datasets. These results demonstrate that our trained models have the potential to be directly employed to real applications.
@@ II Problem formulation
where $n$ stands for the discrete time index. $s(n)$ and $e(n)$ represents the clean source speech and ambient noise, respectively. $a(n)$ denotes RIR and * the convolution operation. In this work, only static speaker is considered, thence the RIR is time-invariant. RIR is composed of the direct-path propagation, early reflections and late reverberation.
We conduct joint speech denoising and dereverberation in this work, which amounts to estimate the (Mel-spectrogram of) desired direct-path speech $x(n)=s(n)*a_{\text{dp}}(n)$ from microphone recording $y(n)$, where $a_{\text{dp}}(n)$ denotes the direct-path part in RIR. The training target of the proposed network and the training signals of neural vocoders will all be derived with $x(n)$.
@@
where $f\in\{0,...,F-1\}$ and $t\in\{1,...,T\}$ denote the indices of frequency and time frame, respectively. $Y(f,t)$, $S(f,t)$ and $E(f,t)$ are the STFT of respective signals, and $X(f,t)$ is the STFT of direct-path speech. $A(f,t)$ is the convolutive transfer function associated to $a(n)$. Convolution * is conducted along time. In the STFT domain, the time domain convolution $s(n)*a(n)$ is decomposed as (frequency-independently) narrow-band convolutions $S(f,t)*A(f,t)$. Speech dereverberation in this work highly relies on learning this narrow-band convolution. For nosie reduction, one important way for discriminating between speech and stationary noise is to test the signal stationarity, which can be modeled in narrow-band as well.
@@ III-A Learning Target: Clean Mel-spectrogram
The power-based or magnitude-based Mel-spectrogram of the target speech $X(f,t)$, denoted as $X_{\text{mel}}(f_{\text{mel}},t)$, can be obtained by weighted summing the squared magnitude $|X(f,t)|^{2}$ or magnitude $|X(f,t)|$ over frequencies with the triangle weight functions of Mel filterbanks, where $f_{\text{mel}}\in\{1,...,F_{\text{mel}}\}$ denotes the index of Mel-frequency. Our preliminary experiments showed that using power-based or magnitude-based Mel-spectrograms achieve similar enhancement performance, thence we can choose either of them according to the Mel setup of ASR or neural vocoder backend models. In this work, we use the power-based Mel-spectrogram according to the setup of commonly-used ASR models. Then, the logMel-spectrogram, namely the logarithm of $X_{\text{mel}}(f_{\text{mel}},t)$
$X_{\text{logmel}}(f_{\text{mel}},t)=\log(\max\{X_{\text{mel}}(f_{\text{mel}},t),\epsilon\})$ (3)
@@
Normally, $\epsilon$ is set to be very small, e.g. 1e-10, to maintain complete speech information. However, in our preliminary experiments, we found that very small speech values are not very informative for both ASR and neural vocoder, and thus can be clipped without harming performance. Moreover, since those small values are highly contaminated by noise or reverberation, the prediction error of them could be very large. For these reasons, we set $\epsilon$ to a relatively large value, e.g. 1e-5 (when the maximum value of time domain signal is normalized. The signal normalization methods will be presented in Section III-D). Fig. 1 gives an example of our target logMel-spectrogram, in which about 40% TF bins are clipped.
@@
$\mathcal{L}_{\text{MAE}}=\frac{1}{F_{\text{mel}}T}\sum_{f_{\text{mel}}=1}^{F_{\text{mel}}}\sum_{t=1}^{T}|\hat{X}_{\text{logmel}}(f_{\text{mel}},t)-X_{\text{logmel}}(f_{\text{mel}},t)|,\end{aligned}$ (4)
where $\hat{X}_{\text{logmel}}(f_{\text{mel}},t)$ is the network output.
@@
$M(f_{\text{mel}},t)=\text{min}\left(\sqrt{\frac{X_{\text{mel}}(f_{\text{mel}},t)}{Y_{\text{mel}}(f_{\text{mel}},t)}},1\right),$ (5)
where $Y_{\text{mel}}(f_{\text{mel}},t)$ denotes the power level of noisy Mel-spectrogram. The square root function transforms the power domain to the magnitude domain. The $\text{min}(\cdot)$ function rectifies the mask into the range of $[0,1]$. The mean squared error (MSE) of the ratio mask is taken as the training loss, namely
$\mathcal{L}_{\text{MRM}}=\frac{1}{F_{\text{mel}}T}\sum_{f_{\text{mel}}=1}^{F_{\text{mel}}}\sum_{t=1}^{T}(M(f_{\text{mel}},t)-\hat{M}(f_{\text{mel}},t))^{2},$ (6)
where $\hat{M}(f_{\text{mel}},t)$ denotes the model prediction of ${M}(f_{\text{mel}},t)$. Then, the enhanced logMel-spectrogram can be obtained as
$\hat{X}_{\text{logmel}}(f_{\text{mel}},t)=\text{log}(\max\{\hat{M}(f_{\text{mel}},t)^{2}Y_{\text{mel}}(f_{\text{mel}},t),\epsilon\}).$ (7)
Figure 1: An example of logMel spectrogram with a clip value of ϵ=italic-$\epsilon=$1e-10 or ϵ=italic-$\epsilon=$1e-5.
@@ III-B The CleanMel Network
Fig. 2 shows the network architecture. The proposed network takes as input (the real $\mathcal{R}(\cdot)$) and imaginary $\mathcal{I}(\cdot)$) parts of) the STFT of microphone recording, i.e. $Y(f,t)$, denoted as $\mathbf{y}$:
@@
where [:]delimited-[]:[:][ : ] represents to take values of one dimension of a tensor. The network is composed of an input layer, interleaved cross-band and narrow-band blocks first in the linear-frequency domain and then in the Mel-frequency domain, a Mel-filterbank, and finally a Linear output layer. The input layer conducts temporal convolution on y with a kernel size of 5, obtaining the hidden representation with the dimensions of $F\times T\times H$. Then one cross-band block and one narrow-band block process the hidden tensor in linear-frequency. Mel-filterbank (with triangle weight functions) transforms the frequency dimension from $F$ linear frequencies to $F_{\text{mel}}$ Mel frequencies, which is realized with a non-trainable matrix multiplication. Then $L$ interleaved cross-band and narrow-band blocks process the tensor in Mel-frequency. After the final narrow-band block, the output Linear layer transforms $H$-dim to 1-dim as either the enhanced logMel-spectrogram or the Mel ratio mask. Note that Sigmoid activation is applied for predicting Mel ratio mask.
@@ III-B1 Narrow-band block
As shown in Eq. (2), the time domain convolution can be decomposed as frequency-independently narrow-band convolutions, while the latter has much smaller complexity compared to the former in terms of the order of room filters. Therefore, modeling the narrow-band convolution would be much more efficient than modeling the time domain convolution. The convolution model of target speech provides not only necessary information for dereverberation of the target speech but also discriminative information between the target speech and other interfering sources. In addition, in narrow-band, non-stationary speech and stationary noise can be well discriminated by testing the signal stationarity. For these reasons, we propose the narrow-band network, which processes frequencies independently along the time dimension, and all frequencies share the same network.
The narrow-band convolution in Eq. (2) is defined between the complex-valued STFT coefficients of source speech and room filter. Thence we process the complex-valued STFT coefficients of noisy signal (in a hidden space), instead of other features such as magnitude, to retain the convolution correlation. We first process in (the finer) linear-frequency with one narrow-band block to fully exploit the convolution information, then after Mel-filterbanks, we process in (the coarser) Mel-frequency with more narrow-band blocks.
The narrow-band network is composed of Mamba blocks [25]. Mamba is a recently proposed network architecture based on structured state space sequence models, which was shown very efficient for learning both short-term and long-term dependencies of sequential data. Besides short-term correlations of signals, there exist some long-term dependencies should be exploited for narrow-band speech enhancement. For example, the convolution model is time invariant in a very long period of time for static speakers. Moreover, Mamba has a linear computational complexity w.r.t time, which is suitable for streaming processing of long audio signals. Specifically, one narrow-band block consists of a forward Mamba for online processing and an optional backward Mamba (averaging with forward output) for offline processing.

### v2

# custom-line = letter = : , command = dashedline , ccommand = cdashedline , tikz = dashed

@@ Abstract
In this work, we propose CleanMel, a single-channel Mel-spectrogram denoising and dereverberation network for improving both speech quality and automatic speech recognition (ASR) performance. The proposed network takes as input the noisy and reverberant microphone recording and predicts the corresponding clean Mel-spectrogram. The enhanced Mel-spectrogram can be either transformed to the speech waveform with a neural vocoder or directly used for ASR. The proposed network is composed of interleaved cross-band and narrow-band processing in the Mel-frequency domain, for learning the full-band spectral pattern and the narrow-band properties of signals, respectively. Compared to linear-frequency domain or time-domain speech enhancement, the key advantage of Mel-spectrogram enhancement is that Mel-frequency presents speech in a more compact way and thus is easier to learn, which will benefit both speech quality and ASR. Experimental results on five English and one Chinese datasets demonstrate a significant improvement in both speech quality and ASR performance achieved by the proposed model. Code and audio examples of our model are available online  [^1] https://audio.westlake.edu.cn/Research/CleanMel.html.
@@ I Introduction
THIS work studies single-channel speech enhancement using deep neural networks (DNNs) to improve both speech quality and automatic speech recognition (ASR) performance. A large class of speech enhancement methods employ DNNs to map from noisy and reverberant speech to corresponding clean speech, conducted either in the time domain [1, 2] or time-frequency domain [3, 4, 5]. While these methods can efficiently suppress noise, they do not necessarily improve ASR performance due to the speech artifacts/distortions introduced by speech enhancement networks [6]. In [7], it was found that time-domain enhancement is more ASR-friendly than frequency-domain enhancement. A time-domain progressive learning method is proposed in [8], which also shows the superiority of time-domain speech enhancement, and the progressive learning mechanism is very effective for robust ASR by mitigating the over-suppression of speech. In [9, 10], ASR performance is largely improved by decoupling front-end enhancement and backend recognition. In [10], it is shown that an advanced time-frequency domain network, i.e., CrossNet, can even outperform the time-domain attentive recurrent network (ARN) [9].
Speech enhancement in Mel-frequency domain, or similarly in rectangular bandwidth (ERB) domain, has been developed under various contexts in the literature. Mel-frequency and ERB bands model human speech perception of spectral envelope and signal periodicity, within which speech enhancement is more perceptually and computationally efficient than within linear-frequency domain or time domain. In [11, 12, 13], spectral envelope enhancement is performed within the ERB bands, and then applying pitch filtering [11, 12] or deep filtering [13] to recover the enhanced speech.
@@
Sub-band networks [3, 14] and full-band/sub-band fusion networks, i.e., FullSubNet [15, 16], have been proposed and achieved outstanding performance. However, separately processing sub-bands in the linear-frequency domain leads to a large computational complexity. To reduce the number of sub-bands and thus the computational complexity, Fast FullSubNet [17] and the work of [18] proposed to perform sub-band processing in the Mel-frequency domain, and then transform back to linear frequency with a joint post-processing network. In [19, 20], speech enhancement is directly conducted in the Mel-frequency domain and then a separately trained neural vocoder is used to recover speech waveform. These methods improve speech enhancement capability by alleviating the burden of enhancing full-band speech details and also by leveraging the powerful full-band speech recovery capacity of advanced neural vocoder, as a result, achieve higher speech quality compared to their linear-frequency counterparts.
@@
In this work, we propose CleanMel, a single-channel Mel-spectrogram enhancement network for improving both speech quality and ASR performance. Different from the previous works [11, 12, 13, 17] that performing speech enhancement in the ERB or Mel-frequency domain, and then applying a joint pitch filtering or deep filtering to obtain the enhanced speech, this work decouples the Mel-spectrogram enhancement and post-processing parts by targeting the enhancement network with clean Mel-spectrogram. The enhanced Mel-spectrogram can be directly used for ASR, or transformed back to waveform with a separate neural vocoder as is done in [19]. Compared to linear-frequency spectrogram or time-domain waveform, Mel-frequency spectrogram presents speech in a more compact and less-detailed way (but still perceptually efficient) and has a lower feature dimension (number of frequencies) from the perspective of machine learning, which would result in lower prediction error. This is beneficial for both speech quality improvement and ASR: (i) Neural vocoders have been extensively studied in the field of Text-to-Speech, and are capable of transforming Mel-spectrogram back to time-domain waveform with sufficient speech details and naturalness. Therefore, the low-error property of Mel-spectrogram enhancement can be hopefully maintained by the neural vocoders and higher speech quality can be achieved. (ii) From the perspective of ASR, there seems no need to first recover the less-accurate full-band (or time-domain sample-wise) speech details and then compress to Mel-frequency. A direct and more-accurate Mel-spectrogram estimation would be preferred.
@@
We adapt the network architecture of our previous proposed (online) SpatialNet [21, 22] with some modifications to better accommodate Mel-spectrogram enhancement. SpatialNet is composed of interleaved cross-band and narrow-band blocks originally proposed for processing multichannel STFT frequencies. The narrow-band block processes STFT frequencies independently to learn the spatial information presented in narrow-band (one frequency), such as the convolutive signal propagation and the spatial correlation of noise. And the cross-band block is designed for learning the across-frequency dependencies of narrow-band information. As for single-channel Mel-spectrogram enhancement in this work, the narrow-band block processes Mel frequencies independently to also learn the (single-channel) convolutive signal propagation of target speech, which is crucial not only for conducting dereverberation of target speech but also for discriminating between target speech and interfering signals. The cross-band block is now reinforced and utilized to learn the full-band spectral pattern in the Mel-frequency domain.
@@
Overall, in this work, we propose a single-channel Mel-spectrogram enhancement network, with an advanced network architecture adapted from SpatialNet [21, 22]. The enhanced Mel-spectrogram can be directly fed into a pre-trained ASR model, or transformed to waveform via a neural vocoder. In addition, we have studied several critical issues when decoupling the Mel-spectrogram enhancement front-end and ASR/Vocoder back-ends. i) We have systematically studied and compared different learning targets that can be used for Mel-spectrogram enhancement, including logMel mapping, Mel ratio masking and the clipping issue of logMel; ii) We have developed a data normalization scheme to align the signal levels of cascaded front-end and back-end models; iii) We have developed an online neural vocoder to enable online speech enhancement.
@@
Experiments are conducted on six public datasets (five English and one Chinese) for speech denoising and dereverberation individually or jointly. Importantly, we adopt a more realistic evaluation setup: from multiple data sources of clean speech, real-measured room impulse responses (RIRs) and noise signals, we collect and organize a relatively large-scale training set, based on which we train the network for once and directly test it on all the six test sets. Experiments show that the proposed model achieves the state-of-the-art (SOTA) speech enhancement performance in term of speech perceptual quality. Moreover, on top of various pre-trained and advanced ASR models, the proposed model prominently improves the ASR performance on all datasets. These results demonstrate that our trained models have the potential to be directly employed to real applications.
@@ II Problem formulation
where $n$ stands for the discrete time index. $s(n)$ and $e(n)$ represents the clean source speech and ambient noise, respectively. $a(n)$ denotes the RIR and * the convolution operation. In this work, only static speaker is considered, thence the RIR is time-invariant. RIR is composed of the direct-path propagation, early reflections and late reverberation.
We conduct joint speech denoising and dereverberation in this work, which amounts to estimate the (Mel-spectrogram of) desired direct-path speech $x(n)=s(n)*a_{\text{dp}}(n)$ from microphone recording $y(n)$, where $a_{\text{dp}}(n)$ denotes the direct-path part in RIR. The training target for both the proposed network and neural vocoders are derived with $x(n)$.
@@
where $f\in\{0,...,F-1\}$ and $t\in\{1,...,T\}$ denote the indices of frequency and time frame, respectively. $Y(f,t)$, $S(f,t)$ and $E(f,t)$ are the STFT of respective signals, and $X(f,t)$ is the STFT of direct-path speech. $A(f,t)$ is the convolutive transfer function associated to $a(n)$. Convolution * is conducted along time. In the STFT domain, the time-domain convolution $s(n)*a(n)$ is approximately decomposed as frequency-wise convolutions $S(f,t)*A(f,t)$. We consider one STFT frequency as one narrow-band, and refer to the frequency-wise convolution as narrow-band convolution. Speech dereverberation in this work highly relies on learning this narrow-band convolution. For noise reduction, one important way for discriminating between speech and stationary noise is to test the signal stationarity, which can be modeled in narrow-band as well.
@@ III-A Learning Target: Clean Mel-spectrogram
[…]
$X_{\text{logmel}}(f,t)=\log(\max\{X_{\text{mel}}(f,t),\epsilon\})$ (3)
@@
Normally, $\epsilon$ is set to be very small, e.g. 1e-10, to maintain complete speech information. However, in our preliminary experiments, we found that very small speech values are not very informative for both ASR and neural vocoder, and can be clipped without harming performance. Moreover, since those small values are highly contaminated by noise or reverberation, the prediction error of them could be very large. For these reasons, we set $\epsilon$ to a relatively large value, e.g. 1e-5 (when the maximum value of time-domain signal is normalized. The signal normalization methods will be presented in Section III-D). Fig. 1 gives an example of our target logMel-spectrogram, in which about 40% TF bins are clipped.
@@
$\mathcal{L}_{\text{MAE}}=\frac{1}{F_{\text{mel}}T}\sum_{f=1}^{F_{\text{mel}}}\sum_{t=1}^{T}|\hat{X}_{\text{logmel}}(f,t)-X_{\text{logmel}}(f,t)|,\end{aligned}$ (4)
where $\hat{X}_{\text{logmel}}(f,t)$ is the network output.
@@
$M(f,t)=\text{min}\left(\sqrt{\frac{X_{\text{mel}}(f,t)}{Y_{\text{mel}}(f,t)}},1\right),$ (5)
[…]
$\mathcal{L}_{\text{MRM}}=\frac{1}{F_{\text{mel}}T}\sum_{f=1}^{F_{\text{mel}}}\sum_{t=1}^{T}(M(f,t)-\hat{M}(f,t))^{2},$ (6)
where $\hat{M}(f,t)$ denotes the model prediction of ${M}(f,t)$. Then, the enhanced logMel-spectrogram can be obtained as
$\hat{X}_{\text{logmel}}(f,t)=\text{log}(\max\{\hat{M}(f,t)^{2}Y_{\text{mel}}(f,t),\epsilon\}).$ (7)
Figure 1: An example of logMel spectrogram with a clip value of $\epsilon=$1e-10 or $\epsilon=$1e-5.
@@ III-B The CleanMel Network
Fig. 2 shows the network architecture. The proposed network takes (the real $\mathcal{R}(\cdot)$) and imaginary $\mathcal{I}(\cdot)$) parts of) the STFT of microphone recording as input, i.e. $Y(f,t)$, denoted as $\mathbf{y}$:
@@
where [:][:][ : ] represents to take values of one dimension of a tensor. The network is composed of an input layer, interleaved cross-band and narrow-band blocks first in the linear-frequency domain and then in Mel-frequency domain, a Mel-filterbank, and a Linear output layer. The input layer performs temporal convolution on y with kernel size 5, producing the hidden representation with the dimensions of $F\times T\times H$. Then one cross-band block and one narrow-band block process the hidden tensor in the linear-frequency domain. A Mel-filterbank (with triangle weight functions) transforms the frequency dimension from $F$ linear frequencies to $F_{\text{mel}}$ Mel frequencies via non-trainable matrix multiplication. Then, $L$ interleaved cross-band and narrow-band blocks process the tensor in Mel-frequency domain.
After the final narrow-band block, the output Linear layer transforms the $H$-dimensional features to 1-dimensional output, producing either the enhanced logMel-spectrogram or the Mel ratio mask. Note that a sigmoid activation is applied when predicting the Mel ratio mask.
@@ III-B1 Narrow-band block
As shown in Eq. (2), the time-domain convolution can be decomposed as frequency-independent narrow-band convolutions, which have much lower complexity than time-domain convolution in terms of room filter order. Therefore, modeling narrow-band convolution is substantially more efficient than modeling time-domain convolution. The convolution model of target speech not only provides necessary information for dereverberation but also helps discriminate target speech from other interfering sources. Additionally, in narrow-band processing, non-stationary speech and stationary noise can be well discriminated by testing the signal stationarity. For these reasons, we propose the narrow-band network, which processes frequencies independently along the time dimension, with all frequencies sharing the same network.
The narrow-band convolution in Eq. (2) is defined between the complex-valued STFT coefficients of source speech and room filters. Therefore, we process the complex-valued STFT coefficients of noisy signal (in a hidden space) rather than other features such as magnitude, to preserve the convolution property. We first process in the finer linear-frequency domain using one narrow-band block to fully exploit the convolution information. After Mel-filterbank transformation, we process in the coarser Mel-frequency domain with additional narrow-band blocks.
The narrow-band network comprises Mamba blocks [25]. Mamba is a recently proposed network architecture based on structured state space sequence models, which has proven highly efficient for learning both short-term and long-term dependencies in sequential data. Besides short-term correlations of signals, there exist some long-term dependencies should be exploited for narrow-band speech enhancement. For example, the narrow-band convolution remains time-invariant over the entire recording for static speakers. Moreover, Mamba has a linear computational complexity w.r.t time, which is suitable for streaming processing of long audio signals. Specifically, each narrow-band block contains a forward Mamba for online processing and an optional backward Mamba (output averaged with forward output) for offline processing.